OpenStudy (bekkah323):

solve: (2m-1)(m+6)/m^3+1 times m^2-m+1/m+6

OpenStudy (bekkah323):

\[\frac{ (2m-1)(m+6) }{ m^3+1 } \times \frac{ m^2-m+1 }{ m+6 }\]

OpenStudy (primeralph):

@Gabriella77 Can handle this. She's smart.

OpenStudy (primeralph):

@Gabriella77 Come one, we need you.

OpenStudy (anonymous):

HA, you know im not good at math. I don't really understand the way the question is written out. @primeralph

OpenStudy (primeralph):

She's smart. She's just tired. Go to bed dear, I'll do it.

OpenStudy (primeralph):

@bekkah323 Solve or simplify?

OpenStudy (bekkah323):

sorry i meant simplify

OpenStudy (skullpatrol):

Are you sure that it is m^3 ?

OpenStudy (bekkah323):

the originial problem is actually \[\frac{ 2m^2+11m-6 }{ m^3+1 } \times \frac{ m^2-m+1 }{ m+6 }\] i was just told to factor the top.

OpenStudy (bekkah323):

nd my book gave me the answer \[\frac{ 2m-1 }{ m+1 }\] i just don't know how they got there

OpenStudy (primeralph):